# Argumentative Experience: Reducing Confirmation Bias on Controversial Issues through LLM-Generated Multi-Persona Debates

### v2

@@ 1. Introduction
Large language models (LLMs) offer a novel design space to explore new ways to support human information seeking. For instance, commercial search engines and services like ChatGPT and Perplexity can synthesize information from multiple sources to provide direct answers to user questions [^1] ChatGPT: https://openai.com/index/chatgpt/, Perplexity: https://www.perplexity.ai/.. Recent work has also explored leveraging the internal knowledge and reasoning capabilities of LLMs to simulate human personas with diverse perspectives (Prpa et al., 2024; Park et al., 2022; Chen et al., 2024; Khan et al., 2024). Imagine a user could generate a set of LLM personas on-demand to debate any topic of interest from multiple perspectives. How should we design such a system to maximize potential benefit to information seekers (e.g., information diversity) while also minimizing associated risks (e.g., LLM hallucination and biases interacting with human biases)?
@@
In traditional information-seeking process, users retrieve facts, evidence, and other information from various sources, often guided by their pre-existing beliefs. This can lead to confirmation bias (Berthet, 2021), where users favor information that supports their prior assumptions. Such biases are prevalent not only in standard information access systems, like search engines, but also in conversational interactions, where users actively shape the direction of the discourse (Sharma et al., 2024; Khan et al., 2024). However, LLM-powered diverse information-seeking could be intentionally designed to provide broad coverage and present contrasting views, regardless of prior user beliefs. Recent design explorations of multi-persona systems have been encouraging, with subjective or indirect user feedback suggesting mitigated confirmation bias (Park et al., 2023a; Chiang et al., 2024; Zhang et al., 2024).
In this work, we explore a novel information seeking experience: letting users launch a debate on any topic of interest in which generated LLM personas present diverse perspectives on the topic. We explore the active and autonomous nature of LLM agents in this setting as they present information and arguments, creating an argumentative experience between humans and AI. In focusing the user experience on exposure to diverse perspectives, we particularly investigated whether exposure to diverse information via generated personas could help reduce user confirmation bias.
@@
Participants in our study launched LLM debates on controversial topics spanning social issues (e.g., ”Should abortion bans be illegal?”) to imaginary scenarios (e.g., ”Should space be colonized?”) We first assessed initial user attitudes toward these topics, then asked users to engage with a baseline search system and our multi-persona debate system. We used a mixed-methods approach to capture participants’ cognitive perceptions, such as real-time eye-tracking metrics, behavioral data, such as user interactions, and qualitative feedback through retrospective think-aloud sessions. These data not only help assess the relative performance of our multi-persona debate system, but also address limitations of other recent exploratory studies on LLM-persona systems, which primarily relied on self-reported metrics and subjective feedback (Park et al., 2023a; Zhang et al., 2024). Our study thus complements this prior work in moving toward a more comprehensive view of how use of LLM personas can impact user information consumption.
Our study investigates three key research questions:
RQ1:
@@
What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
RQ2:
Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
RQ3:
Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
@@
Through a combined analysis of both quantitative and qualitative data, our findings reveal that the argumentative experience powered by LLM personas helps foster constructive deliberation on contrasting viewpoints, making individuals more likely to consume content that challenges their pre-existing beliefs. This happens because persona-based debates present information in a more dynamic, appealing, and persuasive manner, reducing user cognitive load and enhancing interactivity so that people are willing to engage with challenging content. In summary, our work investigates persona-based debate systems to gain deeper insights into human cognitive perceptions and behavioral patterns when interacting with them. We assess the effectiveness of these systems in reducing confirmation bias during information seeking and offer broader design implications to guide the future development of persona-based systems.
@@ 2.1. The Social Simulation and Role-Play of LLM
Over the past five years, LLMs (e.g., GPT, Llama, Gemini, etc.) have rapidly advanced and today represent the state-of-the-art in performance across various natural language tasks [^2] GPT: https://openai.com/index/gpt-4/, Llama: https://www.llama.com/, Gemini: https://gemini.google.com/. Additionally, by being trained on the increasing amounts of web data and computational resources (Kaplan et al., 2020), LLMs have developed strong reasoning and association abilities, as well as internal knowledge that captures social knowledge and cultural norms.
To instruct and tailor content generated by LLMs, prompting is a specific technique used by practitioners, providing initial inputs or cues to a language model (Liu et al., 2023). Researchers have explored creative ways to assess and extend the human-like capabilities of LLMs through prompting (Wang et al., 2021; Wei et al., 2022; Yao et al., 2022). One novel application is the development of agent-based systems where LLMs are used to create highly autonomous agents. LLM personas (or Social Simulacra) are a prominent example. Specifically, by incorporating different beliefs, desires, goals, and characteristics into the prompts, LLMs can role-play to simulate human reactions and behaviors (Shanahan et al., 2023; Park et al., 2022; Chen et al., 2024; Tseng et al., 2024). This approach offers new opportunities for enhancing user experiences in Computer Supported Cooperative Work (CSCW) and other areas of Human-Computer Interaction (HCI). In these fields, generative agents can provide more dynamic, interactive, and autonomous user experiences (Park et al., 2023b). Researchers have explored different use cases, such as using social simulacra to conduct user studies (Argyle et al., 2022; Park et al., 2022), creating generative agents in an interactive sandbox environment (Park et al., 2023b), or directly using LLMs as judges to assist decision-making tasks (Pan et al., 2024), ranging from data annotation (Tan et al., 2024) to language evaluation (Wang et al., 2023).
@@
However, researchers also highlight important challenges in the real-world experiments that employ LLM as the core technique to design persona systems, as well as agent-based systems in general. For example, in particular on the development of social simulacra, LLMs can reflect polarized human opinions regarding different social issues, including misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024). More importantly, the broad idea of agent-based systems also raises important ethical questions about the idea of replacing humans (Prpa et al., 2024). In summary, the development and implementation of LLM-based systems are still in their early stages. Ongoing research is needed to investigate both the advantages and challenges of applying these systems to different real-world scenarios.
2.2. Confirmation Biases in Information Seeking
@@
Confirmation bias is one of the cognitive biases (Berthet, 2021) which refers to the human tendency to seek and interpret evidence that supports their existing beliefs, expectations, or hypothesis (Nickerson, 1998), regardless of the plausibility of those beliefs (Edwards and Smith, 1996). Confirmation bias consists of two aspects: the cognitive aspect (disposition) and the motivational aspect (manifestation) (Berthet et al., 2022). The cognitive aspect of confirmation bias focuses on human tendencies (Berthet et al., 2022; Baron, 2023). It represents a stable psychological disposition that reflects an individual’s tendency toward confirmatory information seeking and evaluation, which can be reliably measured through validated psychometric instruments (Ardi and Pradiri, 2021; Rassin, 2008; Berthet, 2021). The motivational aspect focuses on the relationship with users’ prior beliefs (Baron, 1995). This aspect relates to users’ pre-existing beliefs within specific contexts and manifests dynamically during information-seeking activities, where individuals tend to select information aligning with their attitudes, allocate more attention to confirming content, and assign greater weight to belief-consistent evidence during evaluation (Vedejová and Čavojová, 2022).
@@
Previous research has shown that confirmation bias is widely manifested in seeking behavior, influencing how users interact with information. Specifically, users may be more likely to select and spend more time reading attitude-consistent results (White, 2013; Suzuki and Yamamoto, 2020), while avoiding information that challenges their prior beliefs (El and Moussaoui, 2022; Knobloch-Westerwick et al., 2015; Dickinson, 2020). This behavior may further result in filter bubble and echo chamber (Quattrociocchi, 2017), where like-minded people gather, communicate and interact as an identified community, reinforcing shared beliefs and perspectives (Garrett, 2017).
To mitigate such biases, HCI and CSCW scholars have explored design strategies in visualization and recommendation systems to encourage user engagement with diverse content. One established approach is to increase exposure to opposing viewpoints. For instance, Lex et al. (2018) mixed tweets and hashtags from various partisan groups, Schweiger et al. (2014) used tag clouds to display opposing viewpoints, and Huang et al. (2012) presented counterarguments on a separate screen. Another approach is to increase the system interactivity and user engagement. For instance, Sude et al. (2021) added a ”voting” feature for search results, and Rieger et al. (2021) applied warning labels like ”this search result may reinforce your opinion,” requiring users to click a button to view them. These features help users to be more aware of their information seeking and decision framing process. Additionally, developers created search systems to balance different search results by learning user existing information activities, such as Schwind et al. (2012)’s search system that highlights search result inconsistent with user preferences reflected on their search queries and Jasim et al. (2022)’s bias mitigation recommendation model that suggests diverse sentiment review for online products, informed by user past browsing history. Aforementioned studies suggest that presenting diverse information in a more vivid way can motivate users to engage with content that differs from their prior views.
@@
However, it’s important to recognize that these design and system enhancements remain centered on a static, passive model of information consumption, characteristic of traditional search contexts where information is retrieved and displayed rather than dynamically presented. While increasing exposure to diverse information through external cognitive forcing functions can increase information diversity, this approach may not inherently enhance user engagement or persuasiveness. In contrast, LLM-powered information seeking presents a novel interaction paradigm that enables more diverse, interactive, and dynamic engagement, potentially reducing confirmation bias.
2.3. Opportunities of LLM Constructive Deliberation on Contrasting Viewpoints
@@
Confirmation bias is often associated with polarization effects in which social divides serve to establish and/or reinforce certain viewpoints. While exposure to opposing views in a hostile setting can heighten polarization (Bail et al., 2018), constructive deliberation between opposing views can actually promote empathy, tolerance, and sharing ideas, and reduce polarization (Mutz, 2002; Porter and Schumann, 2018; Levendusky and Stecula, 2021; Bächtiger and Parkinson, 2019; Kim et al., 2021).
Researchers have explored how LLMs can facilitate constructive deliberation and enhance engagement with contrasting viewpoints through mediated discourse.
@@
From the technical side, Tessler et al. (2024) develop the Habermas Machine, a system of LLMs based on Chinchilla which generates statements that express common ground amongst a set of opposing views. Kambhatla et al. (2024) operationalize receptiveness into a computational framework, and use an LLM in combination with the framework to reframe Reddit comments to be more receptive to other users’ comments. Other work also studies grounding, or constructing a shared basis of understanding, in promoting positive discussions on opposing views by fine-tuning specific language model (Cho and May, 2020; Shaikh et al., 2024b).
@@
Researchers in social science and HCI also design LLM-powered tools to facilitate constructive deliberation. For example, Argyle et al. (2023) aim to improve conversation quality in divisive online political conversations using an AI chat assistant. Similarly, Zhang et al. (2024) designed a multi-agent dialogue system in social media conversations. Although no comparison was made with other baseline conditions, they found that participants were influenced either by their own interests or by the system itself to actively explore diverse perspectives, helping to reduce their filter bubbles. Chiang et al. (2024) directly use LLM to play devils’ advocate in AI-assisted group decision-making and found that participants were more likely to form an appropriate trust to adopt AI decisions.
Overall, we see LLMs increasingly being used to reduce polarization and intolerance in ways that might complement LLM-tooling to mitigate related biases. Notably, LLM could be designed to simulate personas that take more critical stances according to predefined roles and characters, which might challenge user prior beliefs. By presenting a new user experience built around diverse feedback and engagements than traditional search, this approach might help counteract the biased information-seeking behaviors commonly seen in traditional search systems.
Research Focus of this Study. Building on prior research, we aim to thoroughly and objectively assess the effectiveness of LLM-generated personas in reducing confirmation bias within an automated debate context. LLM-powered diverse information-seeking via personas demonstrates how LLM agents actively and autonomously can present information and arguments, creating an argumentative experience between humans and AI. Unlike traditional information-seeking that is predominantly user-driven, this argumentative experience could foster a more interactive and collaborative exchange of information to reduce both motivational and cognitive confirmation bias.
While prior studies have examined LLM-generated personas for product recommendations (Park et al., 2023a), reducing filter bubbles through qualitative feedback (Zhang et al., 2024), and generating counterarguments in group decision-making (Chiang et al., 2024), our study objective is to ground the hypotheses (as highlighted above) from a psychological cognitive perspective.
We designed a multi-persona debate system that presents diverse viewpoints through argumentative discourse on controversial issues. By comparing this system with a baseline retrieval-based approach that presents balanced content snippets from opposing perspectives, we evaluate how dynamic, argumentative information presentation influences confirmation bias and belief formation. Through eye-tracking analysis capturing users’ involuntary (bottom-up) and goal-directed (top-down) eye movements, we examine how this active, LLM-driven interaction paradigm shapes user engagement.
In the following sections, we present the system design for the multi-persona debate system (Section 3) and detail the study procedure (Section 4).

### v3

@@ 1. Introduction
Large language models (LLMs) offer a novel design space to explore new ways to support human information-seeking. For instance, commercial search engines and services like ChatGPT and Perplexity can synthesize information from multiple sources to provide direct answers to user questions [^1] ChatGPT: https://openai.com/index/chatgpt/, Perplexity: https://www.perplexity.ai/.. Recent work has also explored leveraging the internal knowledge and reasoning capabilities of LLMs to simulate human personas with diverse perspectives (Prpa et al., 2024; Park et al., 2022; Chen et al., 2024; Khan et al., 2024). Imagine a user could generate a set of LLM personas on demand to debate any topic of interest from multiple perspectives. How should we design such a system to maximize potential benefits to information seekers (e.g., information diversity) while also minimizing associated risks (e.g., LLM hallucination and biases interacting with human biases)?
@@
In the traditional information-seeking process, users retrieve opinions, evidence, and other information from various sources, often guided by their pre-existing beliefs. This can lead to confirmation bias (Berthet, 2021), where users favor information that supports their prior assumptions. Such biases are prevalent not only in standard search engines like Google, but also in newer conversational search systems powered by LLMs, where users can actively steer the discourse, reinforcing the exposure of information that favors their beliefs (Sharma et al., 2024; Khan et al., 2024).
In another line of work (Park et al., 2023a; Chiang et al., 2024; Zhang et al., 2024), scholars explore LLM-enabled multi-persona information-seeking, which presents contrasting viewpoints by simulating different stakeholder perspectives, preventing users from unintentionally seeking evidence that validates their prior beliefs. This design promotes broader and more balanced information coverage with the potential to address confirmation bias as presented in the traditional and newer search systems.
In our study, we further explore this emergent information-seeking experience to understand how it directly reduces confirmation biases. We argue that the user experience offered by multi-personas is particularly argumentative, as simulated personas provide users a constructive debate experience on any topic of interest, presenting diverse perspectives. While recent work has demonstrated the potential of using multi-personas to reduce echo chambers (Zhang et al., 2024) and to encourage critical thinking in group decision-making (Chiang et al., 2024), these studies primarily rely on subjective measurements, such as verbal accounts and user ratings, to assess effectiveness. Although these self-reports provide valuable user insights, they offer limited cognitive evidence for confirming the actual reduction of confirmation bias. To address this, we integrate eye-tracking metrics — supported by established correlations between visual attentional patterns and confirmation biases (detailed in Section 2.2) — as an objective measure to bridge subjective feedback and behavioral cognition.
@@
Participants in our study launched LLM-generated multi-persona debates on controversial topics spanning social issues (e.g., “Should abortion bans be illegal?”) to imaginary scenarios (e.g., “Should space be colonized?”). We first assessed participant attitudes toward eight controversial social topics and then asked them to use a baseline search system and the multi-persona debate system. We used a mixed-methods approach to capture their cognitive perceptions (i.e., real-time eye-tracking metrics), behavioral data (i.e., user interactions), and qualitative feedback through interviews and retrospective think-aloud. These data not only help assess the relative performance of the multi-persona debate system but also address limitations of other recent exploratory studies on LLM-persona systems, which primarily relied on self-reported metrics and subjective feedback (Park et al., 2023a; Zhang et al., 2024). Our study complements this prior work in moving toward a more comprehensive view of how the use of multi-personas can impact user information consumption.
@@
We investigate three key research questions: RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM? RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias? RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence users pre-existing beliefs?
@@
Through a combined analysis of both quantitative and qualitative data, our findings reveal that the argumentative experience powered by LLM personas helps foster constructive deliberation on contrasting viewpoints, making individuals more likely to consume content that challenges their pre-existing beliefs. This happens because persona-based debates present information in a more dynamic, appealing, and persuasive manner, reducing user cognitive load and enhancing interactivity so that people are willing to engage with challenging content. In summary, our work investigates multi-persona debate systems to gain deeper insights into human cognitive perceptions and behavioral patterns when interacting with them. We assess the effectiveness of these systems in reducing confirmation bias during information seeking and offer broader design implications to guide the future development of persona-based systems.
@@ 2.1. The Social Simulation and Role-Play of LLM
LLMs (e.g., GPT, Llama, Gemini, etc.) have rapidly advanced and today represent the state-of-the-art in performance across various natural language tasks [^2] GPT: https://openai.com/index/gpt-4/, Llama: https://www.llama.com/, Gemini: https://gemini.google.com/. By being trained on the increasing amounts of web data and computational resources (Kaplan et al., 2020), these models have developed strong reasoning and association abilities, as well as internal knowledge that captures social knowledge and cultural norms.
To instruct and tailor content generated by LLMs, practitioners commonly use prompting by providing initial inputs or cues to steer the generation (Liu et al., 2023). Researchers have proposed various methods to enhance LLMs’ capabilities through prompting techniques, such as chain-of-thought or interactive reasoning (Wang et al., 2021; Wei et al., 2022; Yao et al., 2022). A notable application in this domain is the creation of agent-based systems. LLM-enabled autonomous agents are designed to simulate complex social behaviors. A prominent example is LLM personas (or Social Simulacra), which enable dynamic, human-like interactions within simulated environments.
In earlier papers that use LLM as social simulacra, Park et al. (2022) embedded traditional UX personas into the prompt, which include “interests (e.g., hobbies, jobs) and personality traits (e.g., kind, bully)”. Their later work (Park et al., 2023b) extends to the incorporation of contextual “summary description (e.g., name, traits, and recent experiences)”. Recent work follows this trend. For example, Park et al. (2023a) use purchase criteria to create assistant agents for camera shopping, Zhang et al. (2024) incorporate character traits “(gender, age, occupation, and education)” to create virtual conversational agents. In related survey paper, Chen et al. (2024) synthesize that persona prompts are dynamically constructed by merging static sociodemographic traits (e.g., age, profession) with emergent behavioral traits shaped through iterative interactions with simulated environments.
Additionally, debate is a key interaction that characterizes how LLM multi-personas engage with one another. This implementation also varies. For example, debates can be structured as simple opposition to user decisions (Chiang et al., 2024), multi-option decision-making scenarios where personas negotiate choices (Park et al., 2023a), or nuanced conversational exchanges based on sociodemographic differences in the persona configuration (Zhang et al., 2024).
The creation of LLM personas offers new opportunities for enhancing user experiences Researchers have explored different use cases, such as using social simulacra to conduct user studies (Argyle et al., 2022; Park et al., 2022), creating generative agents in an interactive sandbox environment (Park et al., 2023b), or directly using LLMs as judges to assist decision-making tasks (Pan et al., 2024), ranging from data annotation (Tan et al., 2024) to language evaluation (Wang et al., 2023). In particular, the debate of LLM multi-personas could be useful in mitigating online polarization and mis/disinformation, one of the core research areas within Computer Supported Cooperative Work & Social Computing (CSCW). We elaborate on this in Section 2.3.
@@
Researchers also highlight challenges in real-world experiments that employ LLM as the core technique to design persona or agent-based systems. For example, on the development of social simulacra, LLMs can reflect polarized human opinions regarding different social issues, including misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024). Additionally, the potential misuse of LLMs could further exacerbate these issues, amplifying existing problems rather than mitigating them (Zhou et al., 2023; Bender et al., 2021). More importantly, the broad idea of agent-based systems also raises important ethical questions about the idea of replacing humans (Prpa et al., 2024). In summary, the development and implementation of LLM-based systems are still in their early stages. Ongoing research is needed to investigate both the advantages and challenges of applying these systems to different real-world scenarios.
2.2. Confirmation Biases and Debiasing Strategies in Information Seeking
2.2.1. The Definition and Measurements of User Confirmation Biases
@@
Confirmation Bias is one of the cognitive biases (Berthet, 2021) which refers to the human tendency to seek and interpret evidence that supports their existing beliefs, expectations, or hypothesis (Nickerson, 1998), regardless of the plausibility of those beliefs (Edwards and Smith, 1996). It consists of two aspects: the Cognitive aspect and the Motivational aspect (Berthet et al., 2022). The cognitive aspect focuses on human tendencies (Berthet et al., 2022; Baron, 2023). It represents a stable psychological disposition that reflects an individual tendency toward confirmatory information seeking and evaluation, which can be reliably measured through validated psychometric instruments (Ardi and Pradiri, 2021; Rassin, 2008; Berthet, 2021). The motivational aspect focuses on the relationship with users’ prior beliefs (Baron, 1995). This aspect contextualizes user beliefs within specific contexts and manifests dynamically during information-seeking activities, such as individuals tend to select information aligning with their attitudes, allocate more attention to confirming content, or assign greater weight to attitude-consistent evidence during evaluation (Vedejová and Čavojová, 2022). Our study measurements cover both aspects to evaluate the reduction of confirmation bias (see details for RQ2 and RQ3 in Section 4.4).
Previous research has explored the influence of confirmation bias on how users seek, browse, select, and engage with attitude-consistent information versus attitude-inconsistent information. In this context, confirmation bias was quantified by comparing either the number or proportion of selected attitude-consistent versus attitude-inconsistent items (Betsch et al., 2001; Jonas et al., 2008; Schulz-Hardt et al., 2000; Ling et al., 2017). For example, Rieger et al. (2021a) calculated the ratio of attitude-consistent to inconsistent selections. Huang et al. (2012a) classified users as demonstrating confirmation bias if more than 50% of their selected items aligned with their prior beliefs.
In addition to methods relying on self-reporting and interaction logs, more unobtrusive methods, such as eye-tracking, were also used in measuring confirmation bias. Eye-tracking is a well-established technique in information retrieval (IR) research (Mostafa and Gwizdka, 2016). It relies on the eye-mind link hypothesis — visual attention follows eye movements (Just and Carpenter, 1980), such as record fixations, scanpaths, pupil dilation, and blinks, to indicate where attention is focused on a screen and how it moves between different areas (Mostafa and Gwizdka, 2016; Siegle et al., 2008). By annotating areas of interest (AOIs) on results, sentences, or words, eye-tracking allows for detailed analysis of how users allocate attention to attitude-consistent versus inconsistent information, revealing confirmation bias (Nelius et al., 2020). Thus, the eye-tracking method was applied to investigate confirmation bias during system interaction in this study (see detailed measurements in Section 4.4)
@@
Previous research has shown that confirmation bias is widely manifested in seeking behavior, influencing how users interact with information when using traditional search engines. Specifically, users may be more likely to select and spend more time reading attitude-consistent results (White, 2013; Suzuki and Yamamoto, 2020), while avoiding information that challenges their prior beliefs (El and Moussaoui, 2022; Knobloch-Westerwick et al., 2015; Dickinson, 2020). This behavior may further result in a filter bubble and echo chamber (Quattrociocchi, 2017), where like-minded people gather, communicate, and interact as an identified community, reinforcing shared beliefs and perspectives (Garrett, 2017).
Advancements in LLM-powered information-seeking have amplified user susceptibility to this bias. For example, Sharma et al. (2024) compared traditional search systems with conversational search powered by LLMs (with and without references). Their results reveal that “participants engaged in more biased information querying” when using conversational search. This trend is consistent with prior research in HCI, CSCW, and IR, which shows that users often rely on heuristics to reduce cognitive dissonance — whether due to conflicting beliefs or a mismatch between beliefs and behaviors, when processing complex information (White and Hassan, 2014; Sparrow et al., 2011; White, 2013). Azzopardi (2021) argues that this information overload exacerbates related biases, including confirmation bias, anchoring bias, and framing effects. Sharma et al. (2024)’s study adds to this growing body of evidence, emphasizing thoughtful design considerations for LLM-powered systems, prioritizing guided interactions over simplistic integration.
2.2.2. Debiasing Strategy in Search Interface Design and System Improvements
A common Debiasing Strategy is to employ design features in visualization and recommendation systems to help users process diverse content more effectively. Though exposing users to diverse perspectives can increase information overload, it remains a crucial first step in mitigating confirmation bias. To address this trade-off, researchers have explored ways to make attitude-inconsistent information more visually appealing, better organized, or concise. For example, Schweiger et al. (2014) used tag clouds to display opposing viewpoints, and Huang et al. (2012b) presented counterarguments on a separate screen to reduce cognitive interference and encourage more balanced consideration.
Another approach is to increase the system interactivity or impose cognitive forcing. For example, Sude et al. (2021) added a “voting” feature for search results, and Rieger et al. (2021b) applied warning labels like “this search result may reinforce your opinion,” requiring users to click a button to view them. These features help users to be more aware of their information-seeking and decision-framing process.
Additionally, developers created search systems to balance different search results by learning user existing information activities, such as Schwind et al. (2012)’s search system that highlights search results inconsistent with user preferences reflected on their search queries and Jasim et al. (2022)’s bias mitigation recommendation model that suggests diverse sentiment review for online products, informed by user past browsing history.
@@ 2.2.2. Debiasing Strategy in Search Interface Design and System Improvements
However, we argue that the aforementioned debiasing strategies remain centered on a static, passive model of information consumption particularly in traditional search experience, where information is retrieved and displayed rather than presented proactively. While increasing exposure through system interactivity or cognitive forcing helps foster user understanding of diverse viewpoints, this may not inherently enhance user engagement or persuasiveness. In contrast, LLM simulation and role-play (described in Section 2.1) presents a novel interaction paradigm that enables more interactive and dynamic engagement, potentially reducing confirmation bias. We introduce this idea along with recent related research in the next section.
2.3. Constructive Deliberation with LLMs
@@
Confirmation bias is often associated with polarization effects in which social divides serve to establish and/or reinforce certain viewpoints. While exposure to opposing views in a hostile setting can heighten polarization (Bail et al., 2018), constructive deliberation between opposing views can actually promote empathy, tolerance, and sharing ideas and reduce polarization (Mutz, 2002; Porter and Schumann, 2018; Levendusky and Stecula, 2021; Bächtiger and Parkinson, 2019; Kim et al., 2021). Scholars in CSCW developed systems to foster constructive deliberation on opposing viewpoints. We found similar design features align closely with those outlined in the previous section.
By employing similar debiasing strategy mentioned in Section 2.2, the ConsiderIt system, designed by Kriplean et al. (2012), “guides people to reflect on trade-offs and the perspectives of others” by presenting pro and con points separately in online debates about U.S. state elections. Systems that are more relevant to crowd information-seeking, such as Poli (Semaan et al., 2015) and PolicyScape (Kim et al., 2019), allow users to gather information reflecting diverse opinions on social media. Additionally, Kim et al. (2021) developed StarryThoughts, a system that visualizes diverse online opinions using the metaphor of “starry nights,” aiming to increase user interactivity.
@@
With the emergence of LLM personas, researchers start applying this to facilitate constructive deliberation. From the technical side, Tessler et al. (2024) developed the Habermas Machine, a system of LLMs based on Chinchilla that generates statements that express common ground amongst a set of opposing views. Kambhatla et al. (2024) operationalize receptiveness into a computational framework and use an LLM in combination with the framework to reframe Reddit comments to be more receptive to other users’ comments. Other work investigates grounding a shared basis of understanding in promoting positive discussions on opposing views by fine-tuning specific language model (Cho and May, 2020; Shaikh et al., 2024b).
@@
Social science and HCI researchers are actively exploring the design of LLM-powered constructive deliberation. For example, Argyle et al. (2023) aims to improve conversation quality in divisive online political conversations using an AI chat assistant. Similarly, Zhang et al. (2024) designed a multi-agent dialogue system in social media conversations. Although no comparison was made with other baseline conditions, they found that participants were influenced either by their own interests or by the system to explore diverse perspectives, helping reduce their filter bubbles actively. Chiang et al. (2024) directly used LLM to play devil’s advocate in AI-assisted group decision-making and found that participants were more likely to form an appropriate trust to adopt AI decisions.
Research Focus of this Study
After reviewing prior work, we argue that LLM multi-personas debate creates an argumentative experience to reduce confirmation bias and polarization. LLM role-play and simulation offer an engaging user experience that is dynamic and proactive (described in Section 2.1). By fostering constructive deliberations (described in Section 2.3), it can mitigate confirmation biases exhibited in traditional and conversational search systems (described in Section 2.2)
Recent work has explored this emergent user experience, as discussed in Section 2.3. However, these studies predominantly validate effectiveness through subjective self-report measures such as verbal accounts and user ratings. While such methods yield valuable user insights, they can be complemented and strengthened by incorporating behavioral and cognitive indicators that directly capture confirmation bias, such as eye-tracking data. Prior research in human perception and psychology has shown a strong positive correlation between visual attention and confirmation bias (as discussed in Section 2.2). Building on this, our work integrates eye-tracking metrics as an objective, real-time measure of attentional patterns, helping to bridge the gap between subjective feedback and quantifiable cognitive behavior.
Specifically, we developed a multi-persona debate system that leverages LLM to foster critical engagement with controversial topics. The system operates by enabling users to create diverse AI personas that debate opposing viewpoints autonomously, encouraging users to synthesize conflicting arguments and mitigate confirmation bias. To evaluate this approach, we compare it against a baseline retrieval-based system that passively displays pre-curated, balanced content snippets from opposing perspectives. Three corresponding research questions are raised. RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM? RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias? RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
Prior work has predominantly focused on RQ1, with qualitative exploration of RQ2 and tangential implications for RQ3. Our study advances this by quantitatively associating confirmation bias reduction with direct belief changes, thereby grounding the efficacy in empirical evidence.